OpenStudy (anonymous):

Which expression is equivalent to 17x − 8x + 2 and why? 2 − 9x, because (17x − 8x) + 2 = −9x + 2 = 2 − 9x 2 + 9x, because (17x − 8x) + 2 = 9x + 2 = 2 + 9x 11x, because 17x − 8x + 2 = 9x + 2 = 11x 7x, because 17x − 8x + 2 = 17x − 10x = 7x

OpenStudy (anonymous):

answer as fast as you can plzand thx

OpenStudy (anonymous):

@aayushi.somani

OpenStudy (ribhu):

2+9X

OpenStudy (ribhu):

2 + 9x, because (17x − 8x) + 2 = 9x + 2 = 2 + 9x